# Simplifying Hypergraph Neural Networks

### v2

@@ 3.2 Simplifying Hypergraph Neural Networks
Linearisation. Recently, some works in graph machine learning (Wu et al., 2019; Wang & Zhang, 2022) show that the message passing mechanism can operate more efficiently without non-linear activation functions while maintaining high effectiveness. Inspired by these works, we further remove the non-linearity from the formulations represented by Eqs. (4$\sim$4d), thereby transforming an L-layer HNN into a single propagation step. For the sake of brevity, we present the simplification results as Lemma 3.1 and provide the mathematical proof for this Lemma in Appendix.
@@
Let $\alpha\in[0,1)$ be a hyperparameter, $\mathbf{X}_{\mathcal{V}}$ be the input node features, and $\hat{\mathbf{X}}_{\mathcal{V}}$ be the output of an $L$-layer UniGCNII/AllDeepSets/ED-HNN/Deep-HGNN. Assume that the learnable parameters and the non-linearity are removed from the model. Then, for any hypergraph =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$, there exists a clique expansion $\mathcal{G}=\{\mathcal{V},\mathbf{W}\}$ to make $\hat{\mathbf{X}}_{\mathcal{V}}=\Big{(}(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}\Big{)}\mathbf{X}_{\mathcal{V}}$.
According to Lemma 3.1, we design the formulation of our SHNN as the following equation:
$\hat{\mathbf{X}}=\mathbf{S}\mathbf{X}_{\mathcal{V}},$ (5)
where $\mathbf{S}=(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$ and we define that the $L$ used to compute $\mathbf{S}$ is the number of layers of an SHNN. To intuitively understand the behaviour of an $L$-layer SHNN, we present the Lemma 3.2. For conciseness, we show the proof of this Lemma in Appendix.
Lemma 3.2.
Let $\mathcal{H}=(\mathcal{V},\mathcal{E},\mathbf{H})$ be a hypergraph and $\mathcal{G}=(\mathcal{V},\mathbf{W})$ be its clique expansion with self-loops. Assume that $\alpha\in[0,1)$, and $\mathbf{S}\!=\!(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$. Then, we have $\mathbf{S}_{ij}>0$ if and only if $v_{i}$ is a $L$-hop neighbour of $v_{j}$ on $\mathcal{H}$, and $\mathbf{S}_{ij}=0$ otherwise.
By Lemma 3.2, an $L$-layer SHNN enables the information exchange between any node $v_{i}$ and its $L$-hop neighbours on the hypergraph. In the next paragraph, we detail the specific designs of $\mathbf{S}$.
Operator design. To apply the SHNN to hypergraph-related downstream tasks, the key is designing $\mathbf{W}$ to generate $\mathbf{S}$. We achieve this using a hyperedge-size-based edge weight that is defined as:
@@
where $\delta(\cdot)$ is a function that returns 1 if $e_{k}$ connect $v_{i}$ and $v_{j}$ and returns 0 otherwise. The main intuition behind the design of this edge weight is that, under certain conditions, the value of $\mathbf{W}_{H_{ij}}$ is positively correlated with the probability of nodes $v_{i}$ and $v_{j}$ having the same label. Consequently, this edge weight design can make the weighted clique expansion fit the homophily assumption, which is that connected nodes tend to be similar to each other (McPherson et al., 2001). We theoretically elaborate on this intuition in Appendix. Furthermore, inspired by previous works in the graph domain (Wu et al., 2020), we generate $\mathbf{S}$ based on a symmetrically normalised $\mathbf{W}_{H}$ with the self-loop. Specifically, we generate the $\mathbf{S}$ by:
@@ 4 Analysis
In this section, we begin by showing the efficiency of the proposed SHNN through a comparison of the training complexity of the SHNN-based framework with the HNN-based framework in hypergraph-related downstream tasks. Additionally, we demonstrate the effectiveness of our SHNN by analyzing its information utilisation and its robustness against the oversmoothing issue.
Training complexity in downstream tasks. According to Eq. (5), the SHNN-based framework for downstream tasks can be summarised as follows:
$\hat{\mathbf{Y}}=\varphi_{\mathbf{\Theta}}(\mathbf{S}\mathbf{X}_{\mathcal{V}}).$ (8)
As illustrated by Eq. (8), the SHNN-based framework for downstream tasks requires training only the task-specific model. Additionally, since the SHNN relies on a pre-defined one-step propagation operator, its computations can be handled during data pre-processing. Consequently, the training complexity of the SHNN-based framework is determined solely by the task-specific model. In contrast, on the basis of Eq. (3), the HNN-based framework involves training both the HNN and the task-specific model. See Figure 1 for an example that compares the training pipelines of the SHNN-based and HNN-based framework. The HNN includes complex learnable operations that facilitate message passing between connected nodes on a hypergraph, inherently making the training of the HNN-based framework more complex than the SHNN-based framework. To provide a precise comparison, we summarize the training complexity of the SHNN-based framework and the four frameworks based on different state-of-the-art HNNs in Table 2. This table shows that the training complexity of the SHNN-based framework is consistently lower than that of the HNN-based frameworks, theoretically confirming the training efficiency of the SHNN-based framework.
Table 2: Training computational complexity of SHNN/HNN-based frameworks. Here, $M$ denote the training computational complexity of the task-specific model, $n$ is the number of nodes, $m$ is the number of hyperedges, $m^{\prime}$ is the number of edges in the clique expansion, $\|\mathbf{H}\|_{0}$ is the number of non-zero entities in $\mathbf{H}$, $L$ is the number of layers, and $d$ is the feature dimension.
SHNN UniGCNII AllDeepSets Deep-HGNN ED-HNN
$\mathcal{O}(M)$ $\mathcal{O}(M+TL(n+m+\|\mathbf{H}\|_{0})d+TLnd^{2})$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$ $\mathcal{O}(M+TLm^{\prime}d+TLnd^{2})$ $\mathcal{O}(M+TL\|\mathbf{H}\|_{0}d+TL(n+m)d^{2})$
Information utilisation. The core strength of existing HNNs lies in their ability to utilize the information embedded within hypergraph structures to generate powerful node features for downstream tasks. Consequently, we analyse the effectiveness of our proposed SHNN by comparing its information utilization capabilities with those of existing HNNs, which is summarised in Proposition 4.1.
@@
Let $H_{0}$ denote the entropy of the information that an HNN utilizes by $L$ message passing function $\phi_{\mathbf{\Theta}}(\cdot)$ introduced in Eq. (1) for generating node features. Similarly, let $H_{1}$ denote the entropy of the information that an $L$-layer SHNN uses for the same purpose. Then, $H_{0}=H_{1}$.
Proof Sketch.
To prove this proposition, we first let $H_{0}^{v_{i}}$ be the entropy of information used by an HNN with $L$ message passing layers in generating the features of node $v_{i}$, $H_{1}^{v_{i}}$ be the entropy of information used by an $L$-layer SHNN for the same purpose, and $H_{2}^{v_{i}}$ denote the entropy of information within node $v_{i}$ and its $L$-hop neighbours. Theron, we prove this proposition by showing that $H_{0}^{v_{i}}=H_{2}^{v_{i}}$ and $H_{1}^{v_{i}}=H_{2}^{v_{i}}$. The detailed proof is presented in Appendix. ∎
@@
Robustness against the oversmoothing issue. According to Lemma 3.2, SHNN can leverage global interactions within the given hypergraph to generate node features by increasing the number of model layers. However, as highlighted in a recent study (Chen et al., 2022), message-passing-based models on hypergraphs may suffer from the oversmoothing issue. This issue refers to the tendency of a model to produce indistinguishable features for nodes in different classes as the model depth increases, which consequently degrades the performance of the learning framework in downstream tasks. Consequently, to further support the effectiveness of SHNN, we theoretically analyse its robustness against the oversmoothing issue. Specifically, we demonstrate the following Proposition:
@@
Let =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$ denote a hypergraph, $\mathcal{G}=\{\mathcal{V},\mathbf{W}_{H}\}$ be its clique expansion with edge weights computed by Eq. (6), $\mathbf{L}$ be the graph Laplacian matrix of $\mathcal{G}$ computed by a symmetrically normalised and self-loops added $\mathbf{W}_{H}$, $\mathbf{X}_{\mathcal{V}}$ represent the input node features, and $\alpha\in[0,1)$ be a hyperparameter. We define that $F(\mathbf{X}^{\prime})=\operatorname{tr}(\mathbf{X}^{\prime\top}\mathbf{L}\mathbf{X}^{\prime})+\frac{\alpha}{1-\alpha}\operatorname{tr}[(\mathbf{X}^{\prime}-\mathbf{X}_{\mathcal{V}})^{\top}(\mathbf{X}^{\prime}-\mathbf{X}_{\mathcal{V}})]$, and $F_{min}$ as the global minimal value of $F(\mathbf{X}^{\prime})$ for $\mathbf{X}\in\mathbb{R}^{n\times d}$. Assume that: 1) $\hat{\mathbf{X}}_{\mathcal{V}}=\mathbf{S}\mathbf{X}_{\mathcal{V}}$; 2) $\mathbf{S}$ is computed by Eq. (7); 3) $\alpha>0$; and 4) $L\to+\infty$. Then, $F_{min}=F(\hat{\mathbf{X}}_{\mathcal{V}})$.
Proof Sketch.
We conduct this proof in two steps. First, we determine the close form for the global minimal point $\mathbf{X}^{\star}$ of $F(\mathbf{X}^{\prime})$. Second, we show that as $L\to+\infty$, the output of an $L$-layer SHNN converges to $\mathbf{X}^{\star}$. We elaborate on this proof in Appendix. ∎
@@
Cora-CA DBLP-CA Citeseer Congress House Senate
# Nodes 2708 41302 3312 1718 1290 282
# Hyperedges 1072 22363 1079 83105 340 315
# Features 1433 1425 3703 100 100 100
# Classes 7 6 6 2 2 2
CE homophily 0.803 0.869 0.893 0.555 0.498 0.509
Node Definition Paper Paper Paper Congressperson Representative Representative
Hyperedge Definition Co-authorship Co-authorship Co-citation Legislative-Bills-Sponsorship Committee Legislative-Bills-Sponsorship
Task NC&HP NC NC & HP NC NC&HP NC
Notably, minimising the first term of $F(\mathbf{X}^{\prime})$ enhances feature similarity among connected nodes on $\mathcal{H}$, and minimising the second term of $F(\mathbf{X}^{\prime})$ encourages the generated features of each node to retain the distinct information from its input features. Therefore, node features that can minimise $F(\mathbf{X}^{\prime})$ capture two key properties: 1) neighbouring nodes on $\mathcal{H}$ have similar features; and 2) each node contains unique information reflecting its individual input characteristics. Consequently, based on Proposition 4.2, an extremely deep SHNN tends to assign similar features to connected nodes on a hypergraph while preserving the unique input information of each node.
According to the discussion above, even if the number of layers in SHNN is increased to extremely large, SHNN can still preserve the unique input information of each node, which enables it to be robust against the oversmoothing issue. This property further ensures the effectiveness of our SHNN. We conduct extensive experiments in Section 6 to support the analysis presented in this section.
5 Related Work
Efficient learning with hypergraphs. Recent efforts have risen to utilize pre-defined hypergraph structures to enhance node feature learning, thereby benefiting downstream tasks (Bick et al., 2023). The predominant models in this area are hypergraph neural networks (HNNs). Inspired by the success of graph neural networks (GNNs) (Wu et al., 2020), most existing works focus on developing HNNs with a message passing mechanism that enables information exchange between connected nodes (Antelmi et al., 2023). Similar to their graph counterparts (Wu et al., 2019; Frasca et al., 2020), HNNs suffer from low learning efficiency. There mainly exist two recent works for this limitation (Feng et al., 2024; Tang et al., 2024). Specifically, (Feng et al., 2024) distills an HNN named HGNN (Feng et al., 2019) into an MLP during the training process, while (Tang et al., 2024) integrates hypergraph structural information into an MLP via a loss function computed by the sum of the maximum node feature distance in each hyperedge. These previous works focus on reducing model inference complexity via a dedicated training process, suitable for scenarios where training and test data are clearly separated. However, many hypergraph-related models are applied in semi-supervised node classification settings that involve the simultaneous usage of training and test data (Feng et al., 2019; Bai et al., 2021; Huang & Yang, 2021; Chien et al., 2022; Wang et al., 2023a, b; Duta et al., 2023; Telyatnikov et al., 2023; Joachims et al., 1999; Chapelle et al., 2009). Hence, reducing the training complexity of hypergraph-related models remains an important yet underexplored question.
We address this challenge by leveraging hypergraph structures via a training-free model. This idea is mainly inspired by a previous work in the graph domain named SIGN (Frasca et al., 2020), which moves the integration of graph information from the training stage to the data preprocessing phase based on training-free graph filters. The main difficulty in implementing this idea for hypergraphs lies in the requirement for hypergraph filters to operate with an incidence tensor (Zhang et al., 2019; Qu et al., 2022), resulting in prohibitively high memory usage for practical applications. Specifically, for a hypergraph with $n$ nodes and a maximum hyperedge size of $k$, the corresponding incidence tensor has $n^{k}$ dimensions. Therefore, instead of using hypergraph filters, we further draw inspiration from SGC (Wu et al., 2019) to introduce a a novel training-free model for hypergraphs, the simplified hypergraph neural network (SHNN). Unlike the SGC approach, which simplifies only the vanilla GCN model (Kipf & Welling, 2017) to create a pure low-pass filter that exclusively smooths the features of connected nodes, our method simplifies four representative Hypergraph Neural Networks (HNNs). This results in a model that not only smooths the features of connected nodes but also preserves the unique input information of each node.

### v3

@@ 3.2.2 Simplifying Hypergraph Neural Networks
Linearisation. Recently, some works in graph machine learning (Wu et al., 2019; Wang & Zhang, 2022) show that the message passing mechanism can operate more efficiently without non-linear activation functions while maintaining high effectiveness. Inspired by these works, we further remove the non-linearity from the functions represented by Eqs. (5$\sim$5d), thereby transforming an MP-Block with $L$ message passing functions into a single propagation step. For the sake of brevity, we present the simplification results as Proposition 3.1 and provide the mathematical proof for this Proposition in Appendix.
Proposition 3.1.
Let $\mathbf{X}_{\mathcal{V}}$ be input node features, $\hat{\mathbf{X}}_{\mathcal{V}}$ be the output of the MP-Block of an UniGCNII/AllDeepSets/ED-HNN/Deep-HGNN with $L$ MP functions, and $\alpha\!\in\![\!0,1\!)$. Assume that learnable parameters and the non-linearity are removed from the block. For any hypergraph =$\mathcal{H}\!=\!\!\{\!\mathcal{V},\!\mathcal{E},\!\mathbf{H}\!\}$, there exists a clique expansion $\mathcal{G}\!=\!\{\!\mathcal{V},\mathbf{W}\!\}$ to make $\hat{\mathbf{X}}_{\mathcal{V}}\!=\!\Big{(}(1\!-\!\alpha)^{L}\!\mathbf{W}^{L}\!+\!\alpha\!\sum_{l=0}^{L-1}(1\!-\!\alpha)^{l}\!\mathbf{W}^{l}\Big{)}\mathbf{X}_{\mathcal{V}}$.
According to Proposition 3.1, we design the formulation of our SMP-Block as the following equation:
$\hat{\mathbf{X}}=\mathbf{S}\mathbf{X}_{\mathcal{V}},$
where $\mathbf{S}=(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$. We define that the $L$ used to compute $\mathbf{S}$ is the number of layers of an SMP-Block. To intuitively understand the behaviour of an $L$-layer SMP-Block, we present the Lemma 3.1. For conciseness, we show the proof of this Lemma in Appendix.
@@
Let $\mathcal{H}=(\mathcal{V},\mathcal{E},\mathbf{H})$ be a hypergraph and $\mathcal{G}=(\mathcal{V},\mathbf{W})$ be its clique expansion with self-loops. Assume that $\alpha\in[0,1)$, and $\mathbf{S}\!=\!(1-\alpha)^{L}\mathbf{W}^{L}+\alpha\sum_{l=0}^{L-1}(1-\alpha)^{l}\mathbf{W}^{l}$. Then, for $i\neq j$, we have $\mathbf{S}_{ij}>0$ if and only if $v_{i}$ is an $L$-hop neighbour of $v_{j}$ on $\mathcal{H}$, and $\mathbf{S}_{ij}=0$ otherwise.
By Lemma 3.1, an $L$-layer SMP-Block enables the information exchange between any node $v_{i}$ and its $L$-hop neighbours on the hypergraph. In the next paragraph, we detail the specific designs of $\mathbf{S}$.
Operator design. To use the SMP-Block in our SHNN, the key is designing $\mathbf{W}$ to generate $\mathbf{S}$. We achieve this using a hyperedge-size-based edge weight that is defined as:
@@
where $\delta(\cdot)$ is a function that returns 1 if $e_{k}$ connects $v_{i}$ and $v_{j}$ and returns 0 otherwise. The main intuition behind the design of this edge weight is that, under certain conditions, the value of $\mathbf{W}_{H_{ij}}$ is positively correlated with the probability of nodes $v_{i}$ and $v_{j}$ having the same label. Consequently, this edge weight design can make the weighted clique expansion fit the homophily assumption, which is that connected nodes tend to be similar to each other (McPherson et al., 2001). We theoretically elaborate on this intuition in Appendix. Furthermore, inspired by previous works in the graph domain (Wu et al., 2020), we generate $\mathbf{S}$ based on a symmetrically normalised $\mathbf{W}_{H}$ with the self-loop. Specifically, we generate the $\mathbf{S}$ by:
@@ 4 Analysis
In this section, we begin by showing the efficiency of the proposed SHNN by comparing its training complexity with the HNNs. Additionally, we demonstrate the effectiveness of our SHNN by analyzing its information utilisation and its robustness against the oversmoothing issue.
Training complexity in downstream tasks. As illustrated by Eq. (1), the SHNN in downstream tasks requires training only the task-specific block. Additionally, since the SHNN relies on a pre-defined one-step propagation operator, its computations can be handled during data pre-processing. Consequently, the training complexity of the SHNN-based framework is determined solely by the task-specific block. In contrast, on the basis of Eq. (2), the HNN involve training both the HNN and the task-specific block. The HNN includes complex learnable operations that facilitate message passing between connected nodes on a hypergraph, inherently making the training of the HNN-based framework more complex than the SHNN. See Figure 1 for an example that qualitatively compares the training pipelines of the SHNN and HNN in node classification. To provide a quantitative comparison, we summarize the training computational complexity of the SHNN and the four state-of-the-art HNNs in Table 1. This table shows that the training complexity of the SHNN is consistently lower than that of the HNN, theoretically confirming the training efficiency of the proposed SHNN.
Information utilisation. The core strength of existing HNNs lies in their ability to utilize the information embedded within hypergraph structures to generate powerful node features for downstream tasks. Consequently, we analyse the effectiveness of our proposed SHNN by comparing its information utilization capabilities with those of existing HNNs and conclude that our methods utilise the same amount of information as exising HNNs without any loss, which is summarised in Proposition 4.1.
@@
Let $H_{0}^{v^{L}_{i}}$ be the entropy of information used by an HNN with $L$ message passing functions in generating the features of a node $v_{i}$, $H_{1}^{v^{L}_{i}}$ be the entropy of information used by an SHNN with an $L$-layer SMP-Block for the same purpose, and $H_{2}^{v^{L}_{i}}$ denote the entropy of information within node $v_{i}$ and its $L$-hop neighbours on the hypergraph. Then, $H_{0}^{v^{L}_{i}}=H_{1}^{v^{L}_{i}}=H_{2}^{v^{L}_{i}}$
@@
Robustness against the oversmoothing issue. According to Lemma 3.1, SHNN can leverage global interactions within the given hypergraph to generate node features by deepening the SMP-Block. However, as highlighted in a recent study (Chen et al., 2022), the message-passing-based models on hypergraphs may suffer from the oversmoothing issue. This issue refers to the tendency of a model to produce indistinguishable features for nodes in different classes as the model depth increases, which consequently degrades the performance of the learning framework in downstream tasks. Consequently, to further support the effectiveness of SHNN, we theoretically analyse its robustness against the oversmoothing issue. Inspired by previous works in the graph machine learning literature (Yang et al., 2021; zhang2020revisiting), we demonstrate the following Proposition:
@@
Let =$\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathbf{H}\}$ denote a hypergraph, $\mathcal{G}=\{\mathcal{V},\mathbf{W}_{H}\}$ be its clique expansion with edge weights computed by Eq. (6), $\mathbf{L}$ be the graph Laplacian matrix of $\mathcal{G}$ computed by a symmetrically normalised and self-loops added $\mathbf{W}_{H}$, $\mathbf{X}_{\mathcal{V}}$ represent the input node features, and $\alpha\in[0,1)$ be a hyperparameter. We define that $F(\mathbf{X}^{\prime})=\operatorname{tr}(\mathbf{X}^{\prime\top}\mathbf{L}\mathbf{X}^{\prime})+\frac{\alpha}{1-\alpha}\operatorname{tr}[(\mathbf{X}^{\prime}-\mathbf{X}_{\mathcal{V}})^{\top}(\mathbf{X}^{\prime}-\mathbf{X}_{\mathcal{V}})]$, and $F_{min}$ as the global minimal value of $F(\mathbf{X}^{\prime})$ for $\mathbf{X}\in\mathbb{R}^{n\times d}$. Assume that: 1) $\hat{\mathbf{X}}_{\mathcal{V}}=\mathbf{S}\mathbf{X}_{\mathcal{V}}$; 2) $\mathbf{S}$ is computed by Eq. (7); 3) $\alpha>0$; and 4) $L\to+\infty$. Then, $F(\hat{\mathbf{X}}_{\mathcal{V}})=F_{min}$.
Notably, minimising the first term of $F(\mathbf{X}^{\prime})$ enhances feature similarity among connected nodes on $\mathcal{H}$, and minimising the second term of $F(\mathbf{X}^{\prime})$ encourages the generated features of each node to retain the distinct information from its input features. Therefore, node features that can minimise $F(\mathbf{X}^{\prime})$ capture two key properties: 1) neighbouring nodes on $\mathcal{H}$ have similar features; and 2) each node contains unique information reflecting its individual input characteristics.
Based on Proposition 4.2, a deep SMP-Block tends to assign similar features to connected nodes on a hypergraph while maintaining each node’s unique input information. Hence, even with an extremely large number of layers, the SMP-Block can preserve the unique input information of each node, making it robust against the oversmoothing issue. This property further ensures the effectiveness of our SHNN. In Section 6, we empirically support the analysis presented in this section.
5 Related Work
Efficient learning with hypergraphs. Recent efforts have risen to utilize existing hypergraphs to enhance node feature learning, thereby benefiting downstream tasks (Bick et al., 2023). The predominant models in this area are hypergraph neural networks (HNNs) with a message passing block that enables information exchange between connected nodes (Antelmi et al., 2023). Similar to their graph counterparts (Wu et al., 2019; Frasca et al., 2020), HNNs suffer from low learning efficiency. There are two recent works for this limitation (Feng et al., 2024; Tang et al., 2024). (Feng et al., 2024) distills an HNN named HGNN (Feng et al., 2019) into an MLP during training, while (Tang et al., 2024) integrates hypergraph structural information into an MLP via a loss function computed by the sum of the maximum node feature distance in each hyperedge. These works focus on reducing model inference complexity via a dedicated training process, suitable for scenarios where training and test data are clearly separated. However, many existing models are applied to semi-supervised hypergraph node classification settings that involve the simultaneous use of training and test data (Feng et al., 2019; Bai et al., 2021; Huang & Yang, 2021; Chien et al., 2022; Wang et al., 2023a, b; Duta et al., 2023; Telyatnikov et al., 2023; Joachims et al., 1999; Chapelle et al., 2009). Therefore, reducing the training complexity of hypergraph-based machine learning models remains an important yet underexplored challenge. We address this challenge with a novel model named simplified hypergraph neural network (SHNN), which is the first to incorporate hypergraph structural information during data preprocessing. This integration significantly improves training efficiency.
@@ 5 Related Work
Name # Nodes # Hyperedges # Features # Classes CE homophily Node Definition Hyperedge Definition Task
Cora-CA 2708 1072 1433 7 0.803 Paper Co-authorship NC&HP
DBLP-CA 41302 22363 1425 6 0.869 Paper Co-authorship NC
Citeseer 3312 1079 3703 6 0.893 Paper Co-citation NC & HP
Congress 1718 83105 100 2 0.555 Congressperson Legislative-Bills-Sponsorship NC
House 1290 340 100 2 0.498 state-of-the-art Committee NC&HP
Senate 282 315 100 2 0.509 state-of-the-art Legislative-Bills-Sponsorship NC